# Echoes and defects in the Calogero model

### v1

@@ Appendix B Solution of Ermakov-Milne equation for the Floquet drive
$\omega^{2}(t)\equiv\omega^{2}_{0}(1+h\cos(\Omega t)),$ (30)
@@
($\mathcal{M}^{S}_{\left(\frac{2{\omega_{0}}}{\Omega}\right)^{2}}\left(\tfrac{1}{2}\Omega t;\left(\frac{2h\omega^{2}_{0}}{\Omega^{2}}\right)\right),\quad\mathcal{M}^{C}_{\left(\frac{4\omega^{2}_{0}}{\Omega^{2}}\right)}\left(\tfrac{1}{2}\Omega t;\left(\frac{2h\omega^{2}_{0}}{\Omega^{2}}\right)\right).$ (31)
@@
where we have used the stationarity of the Wronskian for linearly independent solutions of linear differential equations. We can pick
@@
$\xi^{2}(t)=\left(ax_{1}(t)+bx_{2}(t)\right)^{2}+\frac{1}{4a^{2}}x_{2}(t)^{2},$ (36)
where we used $b_{2}=\frac{1}{2a\mathrm{Wr}(x_{1},x_{2})}=\frac{1}{2a}$. The boundary conditions $\xi(0)=1$ and $\dot{\xi}(0)=0$ follow from the stationarity of the initial state (taken at time $t_{0}=0$) and impose
@@
$\dot{(\xi^{2})}(0)$ $=2\left(ax_{1}(0)+bx_{2}(0)\right)(a\dot{x}_{1}(0)+b\dot{x}_{2}(0))(\tfrac{\Omega}{2})$ (39)
$+\frac{1}{4a^{2}}(2x_{2}(0))\dot{x}_{2}(0)(\tfrac{\Omega}{2})=ab\Omega\equiv 0,$
@@
in which we have omitted for simplicity the time-independent parametric arguments and the dot denotes the time derivative.

### v2

@@ Appendix B Solution of Ermakov-Milne equation for the Floquet drive
$\omega^{2}(t)\equiv\omega^{2}_{0}[1+h\cos(\Omega t)],$ (30)
@@
($\mathcal{M}^{S}_{\left(\frac{2{\omega_{0}}}{\Omega}\right)^{2}}\left[\tfrac{1}{2}\Omega t;\left(\frac{2h\omega^{2}_{0}}{\Omega^{2}}\right)\right],\quad\mathcal{M}^{C}_{\left(\frac{4\omega^{2}_{0}}{\Omega^{2}}\right)}\left[\tfrac{1}{2}\Omega t;\left(\frac{2h\omega^{2}_{0}}{\Omega^{2}}\right)\right].$ (31)
@@
where we use the stationarity of the Wronskian for linearly independent solutions of linear differential equations. We can pick
@@
$\xi^{2}(t)=\left[ax_{1}(t)+bx_{2}(t)\right]^{2}+\frac{1}{4a^{2}}x_{2}(t)^{2},$ (36)
where we used $b_{2}=[{2a\mathrm{Wr}(x_{1},x_{2})}]^{-1}=\frac{1}{2a}$. The boundary conditions $\xi(0)=1$ and $\dot{\xi}(0)=0$ follow from the stationarity of the initial state (taken at time $t_{0}=0$) and impose
@@
$\dot{(\xi^{2})}(0)$ $=2\left[ax_{1}(0)+bx_{2}(0)\right][a\dot{x}_{1}(0)+b\dot{x}_{2}(0)](\tfrac{\Omega}{2})$ (39)
$+\frac{1}{4a^{2}}[2x_{2}(0)]\dot{x}_{2}(0)(\tfrac{\Omega}{2})=ab\Omega\equiv 0,$
@@
in which we omit for simplicity the time-independent parametric arguments and the dot denotes the time derivative.